OpenStudy (anonymous):

how to condense (-1/2)(e^-2 - 1)?

OpenStudy (anonymous):

the answer ended up being (e^2-1)/(2e^2) but i dont know how?

OpenStudy (anonymous):

btw the -2 is the exponent of e

OpenStudy (anonymous):

\[\frac{ -1 }{ 2 } * (\frac{ 1 }{ e^{2}} - 1)\] you can change -1/2 to positive, so long as you change e^2 - 1 to 1 - e^2: \[\frac{ 1 }{ 2 } (1 - \frac{ 1 }{ e^{2} })\] does this help ?